# Shift-cyclicity in Analytic Function Spaces

### v1

@@ 4.2. Multiplier-cyclicity
As $\epsilon>0$ and $g\in\mathcal{H}$ were arbitrarily chosen, we get that $\phi f$ is multiplier-cyclic.
@@
We now list some important properties of the multiplier algebra of a complete Pick space that follow from Theorem 4.3 [[, see]Proposition 2.1]CHMR22.
@@
Let $\mathcal{H}$ be a complete Pick space with kernel $K$, and let $K$, $\mathcal{K}$ and $\mathfrak{b}$ be as in the statement of Theorem 4.3. Then,
@@
For a general RKHS $\mathcal{H}$, not only can we not guarantee that $\mathcal{M}(\mathcal{H})$ is dense, but it may not contain any non-constant functions at all [[, see]Section 2]AM02. Moreover, it is not true for a general RKHS that the kernel functions are multiplier-cyclic [[, see]Theorem 13]Izu17. Thus, Proposition 4.6 provides ways to check whether a given RKHS is not a complete Pick space if it has a normalized irreducible kernel.
@@ 4.2.1. Smirnov representation
We mentioned a result of F. and R. Nevanlinna [NN22] in Section 2.1 (see the references under Theorem 2.3) which states that every $f\in H^{2}(\mathbb{D})$ can be written as $f=\varphi_{1}/\varphi_{2}$ for some $\varphi_{1},\varphi_{2}\in H^{\infty}(\mathbb{D})$. Combining this fact with the Smirnov factorization theorem, it follows that we can choose $\varphi_{2}$ such that it is outer. This is what is called a Smirnov representation of $f$. Taking motivation from this we define the Smirnov class of a RKHS $\mathcal{H}$ as
@@
It is then of interest to ask when
$\mathcal{H}\subset\mathcal{N}^{+}(\mathcal{H})$
holds. First, we note that this is not true in general. For instance, consider $L^{2}_{a}(\mathbb{D})$ and recall that its multiplier algebra is $H^{\infty}(\mathbb{D})$. This means that every function in $\mathcal{N}^{+}(L^{2}_{a}(\mathbb{D}))$ has radial limits a.e. on $\mathbb{T}$ (by Fatou’s theorem [Fat06]). However, it is possible to construct an $L^{2}_{a}(\mathbb{D})$ function that does not have radial limits at any point on $\mathbb{T}$ [[, e.g., use]Theorem 1]Mac62. We therefore conclude that
@@
The following result of Aleman, Hartz, McCarthy and Richter [AHMR17, Theorem 1.1] answers the question in the beginning of this paragraph for complete Pick spaces.
@@
Moreover, $1-\psi$ is multiplier-cyclic for any $\psi\in\mathcal{\mathcal{H}}$ that satisfies (2) above. Thus, $\mathcal{H}\subset\mathcal{N}^{+}(\mathcal{H})$ is always true for a complete Pick space.
Jury and Martin [MJ21, Corollary 3.4] used completely different techniques to obtain the same result but with a different flavor of Smirnov representation.
@@
Let $\mathcal{H}$ be a complete Pick space with a normalized irreducible kernel. Then, for any $f\in\mathcal{H}$ there exist $\varphi,\psi\in\mathcal{M}(\mathcal{H})$ such that $1/\psi\in\mathcal{H}$ and $f=\varphi/\psi$.
It is obvious that if $\varphi_{2}\in\mathcal{M}(\mathcal{H})$ is such that $1/\varphi_{2}\in\mathcal{H}$ then $\varphi_{2}$ is multiplier-cyclic. Thus, Jury and Martin’s result also provides a Smirnov representation for functions in complete Pick spaces.
@@ 2. Inv
We saw that both these properties hold for $H^{2}(\mathbb{D})$, but we do not know the answer for $H^{2}(\mathbb{D}^{2})$ (see Problems 2.192.21). A famous result of Borichev and Hedenmalm [BH97, Theorem 1.4] shows that neither of these properties hold for $L^{2}_{a}(\mathbb{D})$. In fact, they construct an explicit example of a function $f\in L^{2}_{a}(\mathbb{D})$ for which $1/f\in L^{2}_{a}(\mathbb{D})$, but it is not shift-cyclic. They also highlight [[, see]p. 763]BH97 that MC holds for $L^{2}_{a}(\mathbb{D})$, i.e., multiplier-cyclicity is equivalent to shift-cyclicity of the above $f$. The following result by the author settles the case of complete Pick spaces [[, see]Theorem 2.5.4]Sam22.
@@
By the previous discussion, it suffices to show that Mult holds. Thus, let $f_{1},f_{2}\in\mathcal{H}$ be such that $f_{1},f_{2}\in\mathcal{H}$ as well, and let $f_{1}=\varphi_{1}/\psi_{1}$ and $f_{2}=\varphi_{2}/\psi_{2}$ be as in Theorem 4.9. Note that
@@ 5. Cyclicity preserving operators
Up to this point, we were exclusively interested in determining when functions/polynomials in a space are shift/multiplier-cyclic. In cases when we cannot answer with certainty, we were interested in understanding what sort of properties do these functions/polynomials enjoy. In the last part of our discussion, we take an experimental route, and we try to understand other objects that are derived from shift-cyclicity. To be precise, suppose $\mathcal{X}$ and $\mathcal{Y}$ are Banach spaces of analytic functions on some domains (possibly in different dimensions). In this section, we wish to study bounded linear operators $T:\mathcal{X}\to\mathcal{Y}$ that preserve shift-cyclicity, i.e. $Tf$ is shift-cyclic in $\mathcal{Y}$ whenever $f$ is shift-cyclic in $\mathcal{X}$. We refer to such a $T$ as a cyclicity preserving operator. Let us motivate this problem through an example that we already saw.
Recall Theorem 2.15 above, which is a result of Rudin [Rud69, Theorem 4.4.8] that shows that the function
@@
Gibson, Lamoureux and Margrave [GLM11] had a very different motivation for studying cyclicity preserving operators $T:H^{2}(\mathbb{D})\to H^{2}(\mathbb{D})$, coming from a geophysical modelling and signal processing point of view. They noted that all such operators must be weighted composition operators, i.e., there exist holomorphic maps $\varphi:\mathbb{D}\to\mathbb{D}$ and $\psi:\mathbb{D}\to\mathbb{C}$ such that
@@
Mashreghi and Ransford gave a different proof of this result using the Gleason–Kahane–Żelazko theorem [Gle67, KZ68], and generalized it to Banach spaces of analytic functions on $\mathbb{D}$ [[, see]Theorem 3.2]MR15.
@@ 5.1. Maximal domains
Fix $d\in\mathbb{N}$ and let $\mathcal{X}$ be a Banach space of analytic functions on some open set $\Omega\subset\mathbb{C}^{d}$. Recall that $\mathcal{P}_{d}$ is the set of all polynomials in $d$ complex variables. We define the maximal domain of $\mathcal{X}$ [^13] We borrow this terminology from a similar concept related to algebraic consistency. The reader can find more details on this topic in [Har17, MS17]. as
@@
In essence, $\Omega_{max}$ is the largest domain over which point-evaluations $\Lambda_{w}$ make sense for functions in $\mathcal{X}$. The following result of the author [Sam21, Theorem 3.2] provides an important geometric property of $\Omega_{max}$. Recall that the right Harte spectrum $\sigma_{r}(T)$ of a $d$-tuple of commuting operators $T=(T_{1},\dots,T_{d})\in\mathcal{L}(\mathcal{X})$ is a compact set [[, see]Theorem 14.53 (v)]AM02, given by the complement in $\mathbb{C}^{d}$ of
@@
If $\mathcal{X}$ is a Banach space of analytic functions on $\Omega\subset\mathbb{C}^{d}$, then
@@
One usually finds that $\sigma_{r}(S)=\overline{\Omega}$. Then, it is not difficult to determine the maximal domain as one only needs to identify $w\in\partial\Omega$ for which there is an $f\in\mathcal{X}$ with a singularity at $w$. In this case, we conclude that $w\not\in\Omega_{max}$.
Using this idea, it is straightforward to check that the maximal domain of $H^{p}(\mathbb{D}^{d})$ is $\mathbb{D}^{d}$ for all $1\leq p<\infty$ and $d\in\mathbb{N}$. Similarly, for $\mathcal{D}_{t}(\mathbb{D}^{d})$ one can check that
@@
See the recent preprint of Mironov and the author [MS24] for more details and other interesting examples. We also provide the following alternate description of $\Omega_{max}$ and relate it to shift-cyclicity of polynomials [[, see]Theorem 3.1]MS24141414It is important to note that this result holds even if $\mathcal{X}$ is just a topological vector space. We do not need local convexity of $\mathcal{X}$ for this either..
@@
5.2. Main results
@@
$e^{w\cdot z}:=\exp\left(\sum_{j}w_{j}z_{j}\right)\in\mathcal{X}\text{ for all }w\in\mathbb{C}^{d}.$
@@
where $P_{n}$’s are the truncations of the Taylor series of $e^{-w\cdot z}$. Our first main result is the following theorem by the author [Sam21, Theorem 2.1].
@@
$\Lambda(e^{z\cdot w})\neq 0\text{ for all }w\in\mathbb{C}^{d},$
@@
Now, if $T:\mathcal{X}\to\mathcal{Y}$ is a bounded cyclicity preserving operator between two Banach spaces of analytic functions, then it is clear that $\Lambda:=\Lambda_{w}\circ T$ satisfies the hypothesis of the above theorem for each $w$ in the maximal domain of $\mathcal{Y}$. This immediately gives us the first part of the following characterization of cyclicity preserving operators in general [[, see]Theorem 1.5]Sam21.
@@
Moreover, if $\mathcal{Y}=H^{p}(\mathbb{D}^{d^{\prime}})$ for some $1\leq p<\infty$, then every operator as in (5.1) where $\psi$ is shift-cyclic preserves cyclicity.
It is important to note that the second part of the above theorem holds for any $\mathcal{Y}$ that satisfies: (i) the multiplier algebra of $\mathcal{Y}$ is $H^{\infty}(\mathbb{D}^{d^{\prime}})$ and (ii) multiplier-cyclicity is equivalent to shift-cyclicity in $\mathcal{Y}$. However, we do not have a satisfactory answer in general. Therefore, we arrive at the following problem.
@@
Lastly, we note that a result of Chu, Hartz, Mashreghi and Ransford [CHMR22, Theorem 1.2] extends Theorem 5.5 to complete Pick spaces (where maximal domains need to be considered as in [Har17, MS17]). Thus, we close our discussion with one final open problem in this discussion.
@@ Acknowledgements
The author would like to thank Orr Moshe Shalit for motivating him to write this article. Thanks also to Agniva Roy for many helpful suggestions on how to improve the introduction of this article.
[…]

### v3

@@ 4.2. Multiplier-cyclicity
As $\epsilon>0$ and $g\in\mathcal{H}$ were arbitrarily chosen, we get that $\varphi f$ is multiplier-cyclic.
@@
We now list some important properties of the multiplier algebra of a complete Pick space that follow from Theorem 4.3 (see [28, Proposition 2.1]).
@@
Let $\mathcal{H}$ be a complete Pick space with a normalized irreducible kernel $K$, and let $K$, $\mathcal{K}$ and $\mathfrak{b}$ be as in the statement of Theorem 4.3. Then,
@@
For a general RKHS $\mathcal{H}$, not only can we not guarantee that $\mathcal{M}(\mathcal{H})$ is dense, but it may not contain any non-constant functions at all (see [2, Section 2]). Moreover, it is not true for a general RKHS that the kernel functions are multiplier-cyclic (see [55, Theorem 13]). Thus, Proposition 4.6 provides ways to check whether a given RKHS with a normalized irreducible kernel is not a complete Pick space.
@@ 4.2.1. Smirnov representation
We mentioned a result of F. and R. Nevanlinna [79] in Section 2.1 (see the references under Theorem 2.3) which states that every $f\in H^{2}(\mathbb{D})$ can be written as $f=\varphi/\psi$ for some $\varphi,\psi\in H^{\infty}(\mathbb{D})$. Combining this fact with the Smirnov factorization theorem, it follows that we can choose $\psi$ such that it is outer. This is what is called a Smirnov representation of $f$. Taking motivation from this we define the Smirnov class of a RKHS $\mathcal{H}$ as
@@
It is then of interest to ask when we can guarantee the following inclusion
$\mathcal{H}\subset\mathcal{N}^{+}(\mathcal{H}).$
First, we note that this is not true in general. Consider $L^{2}_{a}(\mathbb{D})$ and recall that its multiplier algebra is $H^{\infty}(\mathbb{D})$. This means that every function in $\mathcal{N}^{+}(L^{2}_{a}(\mathbb{D}))$ has radial limits a.e. on $\mathbb{T}$ (by Fatou’s theorem [37]). However, it is possible to construct an $L^{2}_{a}(\mathbb{D})$ function that does not have radial limits at any point on $\mathbb{T}$ (e.g. use [69, Theorem 1]). We therefore conclude that
@@
The following result of Aleman, Hartz, McCarthy and Richter [4, Theorem 1.1] settles the case of complete Pick spaces.
@@
Moreover, $1-\psi$ is multiplier-cyclic for any $\psi\in\mathcal{\mathcal{H}}$ that satisfies (2) and (3) above. Thus, $\mathcal{H}\subset\mathcal{N}^{+}(\mathcal{H})$ is always true for a complete Pick space.
Jury and Martin [57, Corollary 3.4] used completely different techniques to obtain the same result but with a different flavor of Smirnov representation.
@@
Let $\mathcal{H}$ be a complete Pick space with a normalized irreducible kernel. Then, for any $f\in\mathcal{H}$ there exist $\varphi,\psi\in\mathcal{M}(\mathcal{H})$ such that $1/\psi\in\mathcal{H}$ and
$f=\frac{\varphi}{\psi}.$
It is obvious that if $\psi\in\mathcal{M}(\mathcal{H})$ is such that $1/\psi\in\mathcal{H}$ then $\psi$ is multiplier-cyclic. Thus, Jury and Martin’s result also provides a Smirnov representation for functions in complete Pick spaces.
@@ 2. Inv
We saw that both these properties hold for $H^{2}(\mathbb{D})$, but we do not know the answer for $H^{2}(\mathbb{D}^{2})$ (see Problems 2.192.21). A famous result of Borichev and Hedenmalm [23, Theorem 1.4] shows that neither of these properties hold for $L^{2}_{a}(\mathbb{D})$. In fact, they construct an explicit example of a function $f\in L^{2}_{a}(\mathbb{D})$ for which $1/f\in L^{2}_{a}(\mathbb{D})$, but it is not shift-cyclic. They also highlight (see [23, p. 763]) that MC holds for $L^{2}_{a}(\mathbb{D})$, i.e., multiplier-cyclicity is equivalent to shift-cyclicity of the above $f$. The following result by the author settles the case of complete Pick spaces (see [97, Theorem 2.5.4]).
@@
By the previous discussion, it suffices to show that Mult holds. Thus, let $f_{1},f_{2}\in\mathcal{H}$ be such that $f_{1},f_{2}\in\mathcal{H}$ as well, and let
$f_{1}=\frac{\varphi_{1}}{\psi_{1}}\text{ and }f_{2}=\frac{\varphi_{2}}{\psi_{2}}$
be as in Theorem 4.9. Note that
@@ 5. Cyclicity preserving operators
Up to this point, we were exclusively interested in determining when functions/polynomials in a space are shift/multiplier-cyclic. In cases when we cannot answer with certainty, we were interested in understanding what sort of properties do these functions/polynomials enjoy. In the last part of our discussion, we take a slightly unorthodox approach and try to understand other objects that are derived from shift-cyclicity. To be precise, suppose $\mathcal{X}$ and $\mathcal{Y}$ are Banach spaces of analytic functions on some domains (possibly in different dimensions). In this section, we wish to study bounded linear operators $T:\mathcal{X}\to\mathcal{Y}$ that preserve shift-cyclicity, i.e. $Tf$ is shift-cyclic in $\mathcal{Y}$ whenever $f$ is shift-cyclic in $\mathcal{X}$. We refer to such a $T$ as a cyclicity preserving operator. Let us motivate this problem through an example that we already saw.
Recall Theorem 2.15 above, which is a result of Rudin [93, Theorem 4.4.8] that shows that the function
@@
Gibson, Lamoureux and Margrave [40] had a very different motivation for studying cyclicity preserving operators $T:H^{2}(\mathbb{D})\to H^{2}(\mathbb{D})$, coming from a geophysical modelling and signal processing point of view. They noted that all such operators must be weighted composition operators, i.e., there exist holomorphic maps $\varphi:\mathbb{D}\to\mathbb{D}$ and $\psi:\mathbb{D}\to\mathbb{C}$ such that
@@
Mashreghi and Ransford gave a different proof of this result using the Gleason–Kahane–Żelazko theorem [42, 58], and generalized it to Banach spaces of analytic functions on $\mathbb{D}$ (see [72, Theorem 3.2]).
@@ 5.1. Maximal domains
Fix $d\in\mathbb{N}$ and let $\mathcal{X}$ be a Banach space of analytic functions on some open set $\Omega\subset\mathbb{C}^{d}$. Recall that $\mathcal{P}_{d}$ is the set of all polynomials in $d$ complex variables. We define the maximal domain of $\mathcal{X}$ as
@@
We borrow this terminology from a similar concept related to algebraic consistency. The reader can find more details on this topic in [48, Section 5] and [73, Section 2]. In essence, $\Omega_{max}$ is the largest domain over which the point-evaluations $\Lambda_{w}$ make sense for all functions in $\mathcal{X}$. The following result of the author [96, Theorem 3.2] establishes an important geometric property of $\Omega_{max}$. Recall that the right Harte spectrum $\sigma_{r}(T)$ of a $d$-tuple of commuting operators $T=(T_{1},\dots,T_{d})\in\mathcal{L}(\mathcal{X})$ is a compact set (see [2, Theorem 14.53 (v)]), given by the complement in $\mathbb{C}^{d}$ of
@@
If $\mathcal{X}$ is a Banach space of analytic functions on $\Omega\subset\mathbb{C}^{d}$, then [^14] A recent result of Mironov and the author [78] shows that $\Omega_{max}=\sigma_{p}(S^{*})$, i.e, the collection of joint eigenvalues of the $d$-tuple $S^{*}=(S_{1}^{*},\dots,S_{d}^{*})$ of the topological adjoint of $S_{j}$’s.
@@
One usually finds that $\sigma_{r}(S)=\overline{\Omega}$. Then, it is not difficult to determine the maximal domain as one only needs to identify $w\in\partial\Omega$ for which there is an $f\in\mathcal{X}$ with a singularity at $w$. In this case, we conclude that $w\not\in\Omega_{max}$. Using this idea, it is straightforward to check that the maximal domain of $H^{p}(\mathbb{D}^{d})$ is $\mathbb{D}^{d}$ for all $1\leq p<\infty$ and $d\in\mathbb{N}$. Similarly, for $\mathcal{D}_{t}(\mathbb{D}^{d})$ one can check that
@@
See the recent preprint of Mironov and the author [78] for more details and other interesting examples. We also provide the following alternate description of $\Omega_{max}$ and relate it to shift-cyclicity of polynomials (see [78, Theorem 3.1]) [^15] It is interesting to note that this result holds even if $\mathcal{X}$ is just a topological vector space. We do not need local convexity of $\mathcal{X}$ for this either..
@@
5.2. Classification of cyclicity preserving operators
@@
$e^{w\cdot z}:=\exp\left(\sum_{j=1}^{d}w_{j}z_{j}\right)\in\mathcal{X}\text{ for all }w\in\mathbb{C}^{d}.$
@@
where $P_{n}$’s are the truncations of the Taylor series of $e^{-w\cdot z}$. The following theorem by the author [96, Theorem 2.1] is the first step in identifying cyclicity preserving operators.
@@
$\Lambda(e^{w\cdot z})\neq 0\text{ for all }w\in\mathbb{C}^{d},$
@@
Now, if $T:\mathcal{X}\to\mathcal{Y}$ is a bounded cyclicity preserving operator between two Banach spaces of analytic functions, then it is clear that $\Lambda:=\Lambda_{w}\circ T$ satisfies the hypothesis of the above theorem for each $w$ in the maximal domain of $\mathcal{Y}$. This immediately gives us the first part of the following characterization of cyclicity preserving operators in general (see [96, Theorem 1.5]).
@@
Moreover, if $\mathcal{Y}=H^{p}(\mathbb{D}^{d^{\prime}})$ for some $1\leq p<\infty$, then every bounded operator as in (5.1) preserves cyclicity whenever $\psi$ is shift-cyclic.
The second part of the above theorem holds for any $\mathcal{Y}$ that satisfies:
  1. (1)
    the multiplier algebra of $\mathcal{Y}$ is $H^{\infty}(\mathbb{D}^{d^{\prime}})$, and
  2. (2)
    multiplier-cyclicity is equivalent to shift-cyclicity in $\mathcal{Y}$.
However, we do not have a satisfactory answer in general. Therefore, we arrive at the following problem.
@@
Lastly, we note that a result of Chu, Hartz, Mashreghi and Ransford [28, Theorem 1.2] generalizes Theorem 5.5 – in some sense – to complete Pick spaces (where maximal domains need to be considered as in [48, 73]). Thus, we close our discussion with one final open problem.
@@ Acknowledgements
The author would like to thank Orr Moshe Shalit for motivating him to write this article and for providing some important remarks and corrections in the previous draft of the article. Thanks also to Agniva Roy for many helpful suggestions on how to improve the introduction, and to Daniel Seco for his input on the content surrounding Dirichlet-type spaces.